OpenStudy (anonymous):

can somebody help me find the vertex and the x-intercepts of the quadratic equation. y=1/2x^2-4

OpenStudy (anonymous):

To find x-intercepts, put y=0 in the equation and solve for x. Arrange the equation in this way: 2(y+4)=x^2 => vertex = (0,-4)

OpenStudy (anonymous):

thank you so much!!!!!!!!!

OpenStudy (anonymous):

welcome : )